OpenStudy (anonymous):

Prove that (cotX-cscX)(cosX+1)= -sinX

OpenStudy (anonymous):

I worked out the problem and I got it to sinx^2X/sinx= -sinX I just don't know how or where the negative should be

OpenStudy (anonymous):

u shld get a -sin^2 x

OpenStudy (anonymous):

how do you get -sin^2x though

zepdrix (zepdrix):

Oh ok I see where it's coming from, You should be getting something like this, assuming you took the same approach,\[\large \frac{\cos^2x-1}{\sin x}\]Did you get something like that half way through?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

(cosx-1)(cosx+1) = cos^2x - 1 = -sin^2 x

zepdrix (zepdrix):

factor a negative out of each top term :) You have your identity backwards heh!

zepdrix (zepdrix):

\[\large 1-\cos^2x=\sin^2x\]

zepdrix (zepdrix):

\[\large \cos^2x-1=-(1-\cos^2x) \quad = \quad -(\sin^2x)\]

OpenStudy (anonymous):

do you have a program called onenote?

zepdrix (zepdrix):

No :o

zepdrix (zepdrix):

*Googling...* Ooo that looks neat :D

zepdrix (zepdrix):

You still confused on this one? <:o

OpenStudy (anonymous):

haha well i am going to attach a pic of what i have. i just dont know how i get the negative. I understand where you are coming from but I just dont know what the heck i did

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

k

zepdrix (zepdrix):

\[\huge \frac{\cos^2x-1}{\sin x}\quad \color{red}{\neq} \quad \frac{\sin^2x}{\sin x}\]

zepdrix (zepdrix):

You're ok up to that point.

OpenStudy (anonymous):

will you just do the whole problem for me and send me a pic hahah

zepdrix (zepdrix):

Recall your identity,\[\huge \cos^2x+\sin^2x=1\]Moving cosine to the other side gives us,\[\large \huge \sin^2x \quad =\quad \color{purple}{1-\cos^2x}\]

zepdrix (zepdrix):

See the mistake or no? :)

zepdrix (zepdrix):

See how your 1 and cosine are in opposite spots in your step? :D

zepdrix (zepdrix):

That's not necessary, you're only missing one tiny tiny step :D Just reeaddd!! lol

OpenStudy (anonymous):

hold on let me compare what i did to what youre saying!!!

OpenStudy (anonymous):

So my Cos^2X-1/sinx= -sinx is WRONG

zepdrix (zepdrix):

Yes.

OpenStudy (anonymous):

how should it look like

zepdrix (zepdrix):

Because,\[\large \cos^2x-1 \color{red}{\neq} \sin^2x\]

zepdrix (zepdrix):

The top is all we need to worry about.

zepdrix (zepdrix):

Factoring out a negative one gives us,\[\large \cos^2x-1 \quad \rightarrow \quad -(-\cos^2x+1)\]Which we can write like this,\[\large -(1-\cos^2x)\]From here we can apply the identity, \[\large \color{salmon}{1-\cos^2x\quad=\quad \sin^2x}\]Giving us,\[\large -(\color{salmon}{1-\cos^2x}) \quad \rightarrow \quad -(\color{salmon}{\sin^2x})\]

OpenStudy (anonymous):

okayy. let me do this when i get back home around 3. be back online. I MEAN IT

zepdrix (zepdrix):

\[\large \cos^2x-1 \neq \sin^2x\] \[\large 1-\cos^2x=\sin^2x\] Having trouble understanding the difference between those two statements? :C

zepdrix (zepdrix):

lol

OpenStudy (anonymous):

yes I do beucase the one cant be negative

OpenStudy (anonymous):

bye!

zepdrix (zepdrix):

bye :3 go home lol

OpenStudy (anonymous):

okay i am back. but i am trying to work out the problem now

zepdrix (zepdrix):

lol k :)

OpenStudy (anonymous):

hehe i get it now:) THANK YOU

OpenStudy (anonymous):

you were so helpful

zepdrix (zepdrix):

yay team \c:/

OpenStudy (anonymous):

i have another problem i dont know how to do

zepdrix (zepdrix):

Close this thread, Open a new one. You can type @zepdrix in the description somewhere and it will make it easier for me to find it :D

OpenStudy (anonymous):

okay boss